Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?
Answer by stanbon(75887)

Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?
-----------------
EQUATIONS:
R = 2S-4
R+S = 20
-------------
Solve by substitution:
2S-4+S = 20
3S = 24
S = 8 (Sister's age)
--------------
Substitute to solve for R:
R = 2(8)-4
R = 12 (Ricardo's age)
======================================
Cheers,
Stan H.